OpenStudy (anonymous):

What is the angle measure in a scalene triangle with angles 9g, 2g, and 7g?

563blackghost (563blackghost):

Since this is a triangle the angle total would 180 degrees....so we would add the angles in order to make 180 degrees....so this is how our equation would look like... \(\huge{9g+2g+7g=180}\) Combine like terms... \(\huge\color{purple}{18g=180}\) Next you would divide by \(\large{18}\) to get the variable \(\large{g}\) by itself.... \(\Huge\color{red}{\frac{18g}{18}=\frac{180}{18}}\) So what would \(\large{g}\) equal?

OpenStudy (anonymous):

Thank you for the help!

563blackghost (563blackghost):

No problem ^^